OpenStudy (anonymous):

which expression is a cube root of -2?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

A)\[\sqrt[3]{2}(\cos(120°)+isin(120°))\] B)\[\sqrt[3]{2}(\cos(60°)+isin(60°)\]

OpenStudy (anonymous):

C)\[\sqrt[3]{2}(\cos(260°)+isin(260°)\]

OpenStudy (anonymous):

sorry but it is hard

OpenStudy (anonymous):

D)\[\sqrt[3]{2}(\cos(90°)+isin(90°)\]

OpenStudy (anonymous):

D